OpenStudy (barreraa):

Simplify. 5 • 2 • (–4) + 6 ÷ 2 (Points : 4) –37 –23 –17 43

OpenStudy (barreraa):

@AlexandervonHumboldt2

OpenStudy (anonymous):

pemdas

OpenStudy (alexandervonhumboldt2):

5*2*-4+6/2=-40+3=????////